Other things remaining the same, higher is the required return a. lower is the value of...

  1. Other things remaining the same, higher is the required return

    a.

    lower is the value of a bond

    b.

    higher is the value of a stock

    c.

    lower is the value of a stock

    d.

    both a and c

  2. M=1000, n=20, coupon rate = .08, discount rate = .06. The value of the bond is approximately

    a.

    1229.4

    b.

    917.6

    c.

    1129.4

    d.

    none of the above
Answer #1

A) Option d is correct

Higher the required rate of return, will lower the Value of Bond as well as Stock.

Explanation

At 6%,

Value of Bond = $ 1,229.39 (Solved Below)

At 8%,

Value of Bond = $ 999.95

B) Value of the Bond

Value of Bond = Interest * PVIFA(6%,20) + M*PVIF(6%,20)

Interest = 1000*8% = $ 80

Value of Bond = 80 * 11.4699 + 1000 * 0.3118

Value of Bond = $ 1,229.39

Therefore, Option a is correct.